### v2

# To Tag, or Not to Tag: Translating C’s Unions to Rust’s Tagged Unions

Abstract.
Automatic C-to-Rust translation is a promising way to enhance the reliability of legacy system software. However, C2Rust, an industrially developed translator, generates Rust code with unsafe features, undermining the translation’s objective. While researchers have proposed techniques to remove unsafe features in C2Rust-generated code, these efforts have targeted only a limited subset of unsafe features. One important unsafe feature remaining unaddressed is a union, a type consisting of multiple fields sharing the same memory storage. Programmers often place a union with a tag in a struct to record the last-written field, but they can still access wrong fields. In contrast, Rust’s tagged unions combine tags and unions at the language level, ensuring correct value access. In this work, we propose techniques to replace unions with tagged unions during C-to-Rust translation. We develop a static analysis that facilitates such replacement by identifying tag fields and the corresponding tag values. The analysis involves a must-points-to analysis computing struct field values and a heuristic interpreting these results. To enhance efficiency, we adopt intraprocedural function-wise analysis, allowing selective analysis of functions. Our evaluation on 36 real-world C programs shows that the proposed approach is (1) precise, identifying 74 tag fields with no false positives and only five false negatives, (2) mostly correct, with 17 out of 23 programs passing tests post-transformation, and (3) efficient, capable of analyzing and transforming 141k LOC in 4,910 seconds.
Rust, C, Automatic Translation, Union, Tagged Union
journalyear: 2024copyright: rightsretainedconference: 39th IEEE/ACM International Conference on Automated Software Engineering ; October 27-November 1, 2024; Sacramento, CA, USAbooktitle: 39th IEEE/ACM International Conference on Automated Software Engineering (ASE ’24), October 27-November 1, 2024, Sacramento, CA, USAdoi: 10.1145/3691620.3694985isbn: 979-8-4007-1248-7/24/10ccs: Software and its engineering Source code generationccs: Software and its engineering Automated static analysisccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software evolution
1. Introduction
Translating C code to Rust is a promising approach to enhancing the reliability of legacy system software. C Programs often suffer from memory bugs leading to critical security vulnerabilities due to the absence of language-level mechanisms to prevent them (Chen et al., 2011; Thomas, 2019). Rust, a recently developed system programming language, ensures memory safety at compile time through type checking (Matsakis and Klock, 2014; Jung et al., 2017). By translating legacy C code to Rust, developers can detect previously unknown bugs and prevent introducing new bugs (Hutt, 2021).
Since manual translation is laborious and error-prone, an automatic C-to-Rust translator named C2Rust (Wingerter, 2022) has been developed in the industry. It converts C code to Rust by leveraging Unsafe Rust (Rust, 2024g), which allows the use of unsafe language features. These features, such as dereferencing raw pointers and calling functions in external code, are equivalent to C’s features and enable straightforward syntactic translation. However, as the compiler does not ensure their safety, their use contradicts the goal of translation.
To address this, researchers have proposed techniques to reduce the use of unsafe features in C2Rust-generated code by replacing them with safe counterparts in Rust. Laertes (Emre et al., 2021, 2023) and Crown (Zhang et al., 2023) replace raw pointers with references, whose validity is guaranteed by the compiler. Concrat (Hong and Ryu, 2023) replaces certain external function calls by substituting the C lock API with the Rust lock API. Unfortunately, raw pointers and external functions are not the only sources of unsafety, and previous studies have neglected other unsafe features, limiting applicability to the translation of real-world C code.
Unions are an important source of unsafety in C-to-Rust translation that has not been studied yet. A union is a compound data type consisting of multiple fields sharing the same memory storage, facilitating efficient memory use by allowing values of different types to be stored at the same location (Rust, 2024h). Since memory efficiency is crucial in system software, unions are widely used in C. Notably, Emre et al. (Emre et al., 2021) show that 18% of unsafe functions (functions using unsafe features) in C2Rust-generated code involve unions.
Reading a union field is an unsafe feature in Rust because unions do not record which field has been written to. If a program reads a field other than the last-written one, the value is reinterpreted as another type. While reinterpretation is useful for some uses, like packet parsing, it is dangerous in general. For example, reinterpreting an integer as a pointer can lead to invalid memory access. Thus, many C programs avoid reinterpretation when using unions.
To use unions without reinterpretation, it is essential to decide which field to read. Some programs rely on global context to determine the field, but many use tags, i.e., integer values signifying the last-written fields. When using tags, a union and a tag are placed in a single struct, and the program checks the tag before accessing the union’s field. However, tags cannot guarantee the absence of reinterpretation. Programs may read wrong fields after checking tags or set incorrect tag values when writing to fields.
Rust directly supports this pattern of combining tags and unions as a language feature called tagged unions (or enums(Rust, 2024e). This allows defining a tagged union as a single type by enumerating tags and the type of a value associated with each tag. By using tagged unions, programmers can avoid mistakenly reinterpreting values. To access a value in a tagged union, programs must use pattern matching, which checks the tag and provides access to the associated value. When constructing a tagged union, the compiler ensures that the tag and the value’s type match the type definition. Thus, tagged unions are a safe feature in Rust, making it desirable to replace unions accompanied by tags with tagged unions.
C CodeC2RustRust code(unions)Transformer (Section 4)Rust code(tagged unions)Analyzer (Section 3)Analysis result
Figure 1. The workflow of the proposed approach
In this work, we propose techniques to translate C’s unions to Rust’s tagged unions. Fig. 1 shows the workflow of the proposed approach. We first translate C code to Rust code that still contains unions using C2Rust. We then transform the C2Rust-generated code by replacing unions with tagged unions. To enable this transformation, we perform static analysis to obtain information related to unions: (1) the tag field (the field containing a tag value) for each union and (2) tag values associated with each union field. This static analysis must meet several challenging requirements.
First, the analysis needs to determine the values of struct fields. Programs typically use switch/if to access different union fields in different branches, using the tag field in the condition. If a struct field has distinct values when accessing different union fields, it is likely a tag field, and each distinct value is associated with the accessed union field. If it has the same value when accessing different union fields, it is not a tag field. Thus, we can identify tag fields by deciding the value of each struct field in switch/if branches.
To achieve this goal, we propose a must-points-to analysis capable of tracking integer equality. To determine a struct’s field value at each program point based on the branch that the program point belongs to, the struct should be the same as the struct whose field is used in the switch/if condition. Since structs are often passed as pointers, deciding whether they are the same requires must-points-to relations. Additionally, programs sometimes use a local variable storing a field’s value in a condition. In such cases, determining the field’s value in each branch requires the knowledge that the field and the local variable have the same integer.
The second requirement for the analysis is efficiency. The key idea for achieving efficiency is to selectively analyze functions. To identify tag fields, we need the field values only of the structs containing unions. It means that functions not accessing such structs do not need to be analyzed. Therefore, we adopt intraprocedural function-wise analysis, instead of interprocedural whole-program analysis, allowing only the selected functions to be analyzed.
The third requirement is the ability to identify tag values associated with each union field despite the imprecision of the analysis. Given that imprecision is a fundamental limitation of static analysis (Rice, 1953), it may not be possible to determine the field values at some program points. To address this, we propose a heuristic to interpret such partial information. The heuristic involves two steps. First, we examine the accessed union fields and the struct field values in switch/if branches. This provides reliable information because programs typically access the correct union field after checking the tag. However, due to imprecision, it may fail to identify some tag values. Second, we inspect the last-written union fields and the struct field values at each program point. This can capture information missed by the first step but may be incorrect because of an intermediate state, where only the tag or the union field has been set. Therefore, we ignore the field associated with a tag by the second step if it differs from the one associated with the tag by the first step.
Overall, the contributions of this work are as follows:
  • We propose static analysis that identifies tag fields and tag values associated with union fields, consisting of must-points-to analysis capable of tracking integer equality and a heuristic interpreting the analysis results (Section 3).
  • We propose code transformation replacing unions with tagged unions using the analysis results (Section 4).
  • We implement the proposed approach in a tool named Urcrat (union-removing C-to-Rust automatic translator) and evaluate it using 36 real-world C programs. Our evaluation shows that the approach is (1) precise, identifying 74 tag fields with no false positives and only five false negatives, (2) mostly correct, with 17 out of 23 programs passing tests after transformation, and (3) efficient, capable of analyzing and transforming 141k LOC in 4,910 seconds (Section 5).
We also discuss related work (Section 6) and conclude the paper (Section 7).
2. Background
In this section, we briefly describe the use of unions with tags in C (Section 2.1), how C2Rust translates such C code to Rust (Section 2.2), and how tagged unions in Rust can safely represent the same logic (Section 2.3).
2.1. Unions with Tags
As an example of C code using unions, we use the syntax and evaluation of simple arithmetic expressions defined as follows:
[…]
An expression is either a constant $1$, a negation, an addition, or a multiplication. This syntax is implemented as follows:
struct Expr {
int kind; union { struct Expr *e; struct BExpr b; } v;
struct BExpr { struct Expr *l; struct Expr *r; };
struct Expr is the type of an expression, and its field kind indicates the kind of expression: 0 for constant $1$, 1 for negation, 2 for addition, and 3 for multiplication. The inner union value v stores the necessary data for each kind of expression. When kind is 1, the operand of negation is stored in v.e; when kind is 2 or 3, the operands are stored in v.b.l and v.b.r. Since kind signifies which union field has been written to, it is the tag field for the union.
A function evaluating an expression is implemented as follows:
int eval(struct Expr *e) {
switch (e->kind) {
case 0: return 1;
case 1: return -eval(e->v.e);
case 2: return eval(e->v.b.l) + eval(e->v.b.r);
case 3: return eval(e->v.b.l) * eval(e->v.b.r);
default: exit(1); }}
It evaluates the expression by checking the kind field and accessing the appropriate union field accordingly. When kind is 1, it accesses the union field e; when kind is 2 or 3, it accesses b. No field is accessed otherwise.
Programs sometimes use if to check tags, particularly to compare with a specific tag value. An example using if is shown below:
if (e->kind == 1) return -eval(e->v.e);
While tag fields are beneficial for accessing the correct union fields, they cannot prevent memory bugs. For example, the following code accesses e->v.b despite kind being 1:
switch (e->kind) {
case 1: return eval(e->v.b.l) + eval(e->v.b.r);
Here, e->v.b.l accesses the pointer stored in e->v.e, but e->v.b.r reads an arbitrary value, potentially causing invalid memory access. Furthermore, even when eval is correctly implemented, an incorrect tag value can be assigned during the construction of an Expr. The following code sets e.kind to 2 but writes to e.v.e:
struct Expr e; e.kind = 2; e.v.e = ...;
Passing a pointer to e to eval results in invalid memory access.
2.2. C2Rust’s Translation
C2Rust translates the definition of struct Expr to Rust as follows:
struct Expr { kind: i32, v: C2RustUnnamed }
union C2RustUnnamed { e: *mut Expr, b: BExpr }
struct BExpr { l: *mut Expr, r: *mut Expr }
Since Rust does not support anonymous types, C2Rust gives the name C2RustUnnamed to the union. If a file contains multiple anonymous types, they get C2RustUnnamed_$n$ where $n$ is a unique integer.
The function eval is translated as follows:
fn eval(e: *mut Expr) -> i32 {
match (*e).kind {
0 => return 1,
1 => return -eval((*e).v.e),
2 => return eval((*e).v.b.l) + eval((*e).v.b.r),
3 => return eval((*e).v.b.l) * eval((*e).v.b.r),
_ => exit(1), }}
Since Rust provides match statements instead of switch, the function employs match to check the tag. While match is mainly used for pattern matching on tagged unions, it can also handle integers, similar to switch, but without fall-through behavior.
2.3. Tagged Unions
We can implement the same syntax using tagged unions as follows:
struct Expr { v: C2RustUnnamed }
enum C2RustUnnamed {
One, Neg(*mut Expr), Add(BExpr), Mul(BExpr) }
struct BExpr { l: *mut Expr, r: *mut Expr }
In Rust, the enum keyword defines tagged unions. Although the name C2RustUnnamed is impractical, we retain it for consistency with the C code. A tagged union’s definition lists its variants, i.e., values with distinct tags. Each tag is an identifier, not an integer, and the type of a value associated with each tag is specified after the tag. The defined tagged union has four variants with tags One, Neg, Add, and Mul. The Neg tag is associated with an Expr pointer, while Add and Mul are associated with a BExpr value. Since C2RustUnnamed now contains the tag, the kind field in the struct is no longer necessary.
Now, eval can be implemented with pattern matching as follows:
1fn eval(e: *mut Expr) -> i32 {
2 match (*e).v {
3 C2RustUnnamed::One => return 1,
4 C2RustUnnamed::Neg(e) => return -eval(e),
5 C2RustUnnamed::Add(b) => return eval(b.l)+eval(b.r),
6 C2RustUnnamed::Mul(b) => return eval(b.l)*eval(b.r),
7 }}
Each pattern matching branch specifies a tag and binds the associated value to an identifier. For instance, in line 4, the associated value is bound to e and then passed to eval. Since the compiler ensures all variants are covered, the _ (default) branch is unnecessary.
The code pattern using if to check tags is also supported through if-let (Rust, 2024d), which is another form of pattern matching. The following code performs computation only when the tag is Neg:
if let C2RustUnnamed::Neg(e) = (*e).v {
return -eval(e); }
Pattern matching enables the compiler to detect programmers’ mistakes that can cause memory bugs. Consider the following Rust code where the value associated with Neg is treated as a BExpr type:
match (*e).v {
C2RustUnnamed::Neg(b) => return eval(b.l) + eval(b.r),
The type of b is *mut Expr, as specified in the type definition. Since *mut Expr lacks the fields l and r, the compiler raises an error.
Additionally, the compiler verifies the correct construction of tagged union values. Consider the following code, which incorrectly initializes a tagged union with the tag Add and an Expr pointer:
let e1: Expr = ...;
let e2 = Expr { v: C2RustUnnamed::Add(&mut e1) };
Since the type definition requires a BExpr value for Add, the code does not pass type checking.
3. Static Analysis
In this section, we present static analysis designed to facilitate the transformation of unions with tags into tagged unions. The objectives of this analysis are to (1) identify a tag field for a union, if one exists, and (2) determine the tag values associated with each union field. The proposed static analysis consists of four steps:
  1. (1)
    Identification of candidate structs, those containing unions and their potential tag fields (Section 3.1).
  2. (2)
    Whole-program may-points-to analysis (Section 3.2).
  3. (3)
    Intraprocedural must-points-to analysis for selected functions (Section 3.3).
  4. (4)
    Interpretation of the analysis results using a heuristic (Section 3.4).
As we analyze Rust code generated by C2Rust, rather than the original C code, code examples in this section are written in Rust.
3.1. Candidate Identification
The first step of the analysis is to identify structs that likely contain unions and their tag fields. We also determine which functions to analyze based on the identified structs. If no such structs are found, the analysis terminates at this step, concluding that the program does not contain unions to be transformed into tagged unions.
To concretely define candidate structs, we first define tag-eligible fields, those that can potentially serve as tag fields. A field of a struct is considered tag-eligible if (1) it has an integer type, (2) when it appears on the left side of an assignment, the operator is =, excluding others such as +=, and (3) it is never referenced by a pointer. For the first condition, integer types include bool, i8, u8, i16, u16, i32, u32, i64, and u64, which C2Rust translates from _Bool, signed/unsigned char, short, int, and long. The second and third conditions arise from limitations in the expressibility of tagged unions. Tags of tagged unions are not integers and thus cannot undergo integer operations. Moreover, since tags do not exist as fields, they cannot be referenced. Consequently, fields that exhibit such behavior cannot be transformed into tagged unions.
We also define candidate unions, which can potentially be accompanied by tags. A union is a candidate if (1) it is a field of a struct that contains at least one tag-eligible field, and (2) its name begins with C2RustUnnamed. The second condition indicates that the union is anonymous in the C code. If a union has a name, it can be used independently of the struct, not being expressible as tagged unions.
We finally define candidate structs. A candidate struct is a struct containing at least one candidate union.
We now describe how to determine which functions to analyze. To identify tag fields for unions and values associated with union fields, we need to ascertain the possible values of candidate struct fields. If a function neither reads from nor writes to a field, intraprocedural analysis of the function provides no information about that field’s value. Therefore, we analyze only the functions that access fields of candidate structs. Such functions can be easily identified through syntactic examination.
3.2. May-Points-To Analysis
The second step of the analysis is to conduct a whole-program may-points-to analysis. This step is essential because may-points-to relations are required to ensure the soundness of the subsequent must-points-to analysis. The utilization of may-points-to relations in the must-points-to analysis is detailed in Section 3.3.
We employ a field-sensitive Andersen-style analysis (Pearce et al., 2007) for the may-points-to analysis. This analysis is flow-insensitive and has a time complexity of $O(n^{3})$. Although other may-points-to analyses exist, such as field-insensitive Andersen-style (Andersen, 1994) and Steensgaard-style analyses (Steensgaard, 1996b, a), they are too imprecise, leading to unacceptably imprecise must-points-to analysis results.
3.3. Must-Points-To Analysis
The third step is to perform an intraprocedural must-points-to analysis for each selected function. Our overall algorithm is akin to typical must-points-to analyses (Kastrinis et al., 2018). The execution state at each program point is represented as a graph, with nodes denoting memory locations and edges expressing points-to relations. The analysis iteratively updates the state at each program point until reaching a fixed point. Each state is derived by joining the state from the previous iteration with the state resulting from applying the current instruction’s effect to the previous program point’s state.
We first describe how we visualize graphs throughout this section. Since nodes denote memory locations, some nodes correspond to the stack locations used by local variables. For clarity in visualization, we draw a dashed arrow from the name of a local variable to the node representing its memory location. Note that a variable name is not a node and thus this arrow is not an edge. For example, x = &y constructs the following graph:
xy
The graph has two nodes and one edge and indicates that the pointer at the memory location of x points to the memory location of y.
Since our goal is not only to compute must-points-to relations but also to track fields’ integer values, we optionally label each node with @$N$, where $N$ is a set of integers. The label @$N$ encodes integer value information in pointer graphs. If a node is unlabeled, the value at the location is a usual C value (such as an integer, pointer, struct, union, or array). If a node is labeled @$N$, the value at the location is an imaginary value at addresses $N$, i.e., the possible addresses of the location are $N$. Consequently, if a node $v$ has an outgoing edge to a node labeled @$N$, then the possible values at $v$’s location are $N$. For instance, x = 1 constructs the following graph:
x @1
Using the @$N$ label is beneficial since it enables efficient propagation of integer values to memory locations known to hold the same value. Consider the following graph, constructed by x = y:
xy
If we obtain the fact that x equals 1, we update the graph as follows:
xy @1
Then, we automatically discover that y also equals 1. As this example demonstrates, the @$N$ label facilitates the update of the value at multiple memory locations by labeling only a single node.
We now discuss how to analyze code involving unions. Consider the following example, where Expr is defined as in Section 2.2:
1fn eval(e: *mut Expr) -> i32 {
2 let k = (*e).kind;
3 match k {
4 1 => return -eval((*e).v.e),
After line 2, we have the following graph:
ek.kind
Here, the edge labeled .kind indicates the presence of a struct/union at the location of the node where the edge originates, with the pointer stored in the kind field referring to the pointed node’s location. Since line 3 uses k as the condition for match, we determine that k equals 1 in line 4. Consequently, the graph at the beginning of line 4 is as follows:
ek @1 .kind
From this graph, we conclude that the value of kind is 1 when the union field e is accessed in line 4.
More complex scenarios involve joining two graphs. Consider the following code example:
1match (*e).kind {
2 2 => ...,
3 3 => ...,
4 _ => return, }
5let lv = eval((*e).v.b.l);
In lines 2 and 3, we have the following graphs, respectively:
e @2 .kinde @3 .kind
Since line 5 is reachable from both lines 2 and 3, we need to join the graphs. When joining graphs, the integer sets in the labels are unioned, resulting in the following graph:
e @2,3 .kind
From the graph, we conclude that the value of kind is 2 or 3 when the union field b is accessed in line 5.
The join of graphs is carefully defined to maintain the soundness of the analysis. Consider the following code, where control flow splits based on the value of kind and then merges:
1if (*e).kind == 1 { ... }
2else { ... }
3...
The states in lines 1 and 2 are as follows, respectively:
e @1 .kinde.kind
In line 1, (*e).kind is known to be 1. However, in line 2, its value is unknown. When entering line 3, these graphs are joined as follows:
e.kind
[…]
Input : $g_{1}$, $g_{2}$
Output : $g$
1 $g.\mathsf{nodes}:=\emptyset$$g.\mathsf{edges}:=\emptyset$$\mathsf{worklist}:=\emptyset$;
[…]
3      if $g_{1}\ \text{has a node}\ v_{1}\ \text{corresponding to}\ x$:
4           if $g_{2}\ \text{has a node}\ v_{2}\ \text{corresponding to}\ x$:
5                $g.\mathsf{nodes}.\mathsf{insert}((v_{1},v_{2}))$;
6                $\mathsf{worklist}.\mathsf{insert}((v_{1},v_{2}))$;
7
8
9
10                while $\mathsf{worklist}\not=\emptyset$:
11                     $(v_{1},v_{2}):=\mathsf{worklist}.\mathsf{pop}()$;
12                     for $(v_{1},v_{1}^{\prime},f)\leftarrow v_{1}\text{'s outgoing edges in}\ g_{1}$:
13                          if $g_{2}\ \text{has an edge}\ (v_{2},v_{2}^{\prime},f)$:
14                               $g.\mathsf{edges}.\mathsf{insert}((v_{1},v_{2}),(v_{1}^{\prime},v_{2}^{\prime}),f)$;
15                               if $(v_{1}^{\prime},v_{2}^{\prime})\not\in g.\mathsf{nodes}$:
16                                    $g.\mathsf{nodes}.\mathsf{insert}((v_{1}^{\prime},v_{2}^{\prime}))$;
17                                    $\mathsf{worklist}.\mathsf{insert}((v_{1}^{\prime},v_{2}^{\prime}))$;
18
19
20
21                                    if $v_{1}\ \text{has a label}\ {\sf@}N_{1}\ \text{in}\ g_{1}$:
22                                         if $v_{2}\ \text{has a label}\ {\sf@}N_{2}\ \text{in}\ g_{2}$:
23                                              $\text{set}\ (v_{1},v_{2})\text{'s label to}\ {\sf@}(N_{1}\cup N_{2})\ \text{in}\ g$;
24
25
26
Algorithm 1 Graph joining
Algorithm 1 illustrates the algorithm for graph joining. Edges are intersected to retain must-point-to relations, while integer sets in labels are unioned. In the pseudo-code, each node in the input graphs $g_{1}$ and $g_{2}$ is denoted by a unique identifier $v$, and each node in the output graph $g$ is represented by a pair of identifiers $(v_{1},v_{2})$. Each edge is represented by $(v,v^{\prime},f)$, denoting an edge labeled $f$ from node $v$ to node $v^{\prime}$. Unlabeled edges are treated as edges with the empty label $\epsilon$. Initially, $g$ and the worklist, which stores $g$’s nodes to be visited, are both empty (line 1). Then, we add nodes corresponding to local variables to $g$ (lines 2–6). Finally, we visit each added node until the worklist is empty (lines 7–8). During the visit, edges with the same label in $g_{1}$ and $g_{2}$ are added to $g$ (lines 9–14), and node labels are unioned if they exist (lines 15–17).
We now discuss the utilization of may-points-to relations during the must-points-to analysis. Consider the following code:
1let e = if ... { &mut ev } else { ... };
2let k = (*e).kind;
3ev = ...;
4match k {
5 1 => ...,
Since e may not point to ev, the state after line 1 is as follows:
eev
This graph indicates that e points to some location, not necessarily the same as ev’s location. Line 2 updates the graph as follows:
ek.kindev
Line 3 mutates the value of ev, possibly changing the value of (*e).kind. Thus, (*e).kind is not necessarily equal to k after line 3, necessitating an appropriate update to the graph. As the graph itself does not reveal any relations between e and ev, we rely on the precomputed may-points-to relations, which indicate that e may point to ev. Consequently, we remove all outgoing edges, including the kind edge, from *e’s node, resulting in the following graph:
ekev
In line 5, we successfully avoid the incorrect conclusion that (*e).kind equals 1, as shown in the following graph:
ek @1 ev
Like this, the analysis removes the appropriate edges from the graph at each (indirect) assignment and function call according to the may-points-to relations. The effect of a function call is equivalent to the cumulative effects of all assignments reachable by the call.
3.4. Analysis Result Interpretation
1 def IdentifyTags(struct $s$, union $u$, field $f_{s}$):
2      $res:=$ CollectFromAccesses($s,u,f_{s}$);
3      if $res=\mathsf{None}$:
4           return $\mathsf{None}$;
5
6           $(\mathsf{field\_tags},\mathsf{access\_tags}):=res$;
7           $\mathsf{field\_tags}^{\prime}:=$ CollectFromStructs($s,u,f_{s}$);
8           $\mathsf{struct\_tags}:=\emptyset$;
9           for $f_{u}\leftarrow\text{fields of}\ u$:
10                $\mathsf{tags}:=\mathsf{field\_tags^{\prime}}[f_{u}]\setminus\mathsf{access\_tags}$;
11                if $\mathsf{struct\_tags}\cap\mathsf{tags}\not=\emptyset$:
12                     return $\mathsf{None}$;
13
14                     $\mathsf{field\_tags}[f_{u}]:=\mathsf{field\_tags}[f_{u}]\cup\mathsf{tags}$;
15                     $\mathsf{struct\_tags}:=\mathsf{struct\_tags}\cup\mathsf{tags}$;
16
17                     $\mathsf{all\_tags}:=$ CollectAllTags($s,u,f_{s}$);
18                     $\mathsf{rem\_tags}:=\mathsf{all\_tags}\setminus(\mathsf{access\_tags}\cup\mathsf{struct\_tags})$;
19                     return $(\mathsf{field\_tags},\mathsf{rem\_tags})$;
20
21                     def CollectFromAccesses(struct $s$, union $u$, field $f_{s}$):
22                          $\mathsf{field\_tags}:=\mathsf{Map}()$; $\mathsf{all\_tags}:=\emptyset$;
23                          for $l\leftarrow\text{analyzed program points}$:
24                               if a $\text{a field}\ f_{u}\ \text{of}\ u\ \text{is accessed}\ \text{at}\ l$:
25                                    $N:=\text{possible values of}\ f_{s}\ \text{at}\ l$;
26                                    if $N\ \text{is from}\ \texttt{\small if}\ or\ \texttt{\small match}$:
27                                         $\mathsf{tags}:=N\setminus\mathsf{field\_tags}[f_{u}]$;
28                                         if $\mathsf{all\_tags}\cap\mathsf{tags}\not=\emptyset$:
29                                              return $\mathsf{None}$;
30
31                                              $\mathsf{field\_tags}[f_{u}]:=\mathsf{field\_tags}[f_{u}]\cup\mathsf{tags}$;
32                                              $\mathsf{all\_tags}:=\mathsf{all\_tags}\cup\mathsf{tags}$;
33
34
35
36                                              return $(\mathsf{field\_tags},\mathsf{all\_tags})$;
37
38                                              def CollectFromStructs(struct $s$, union $u$, field $f_{s}$):
39                                                   $\mathsf{field\_tags}:=\mathsf{Map}()$;
40                                                   for $l\leftarrow\text{analyzed program points}$:
41                                                        if $l\ \text{is end of a basic block}$:
42                                                             for $v\leftarrow\text{struct}\ s\ \text{reachable at}\ l$:
43                                                                  if $\text{union}\ u\ \text{in}\ v\ \text{has a field}\ f_{u}$:
44                                                                       $N:=\text{possible values of}\ f_{s}\ \text{of}\ v$;
45                                                                       $\mathsf{field\_tags}[f_{u}]:=\mathsf{field\_tags}[f_{u}]\cup N$;
46
47
48
49
50                                                                       return $\mathsf{field\_tags}$;
51
52                                                                       def CollectAllTags(struct $s$, union $u$, field $f_{s}$):
53                                                                            $\mathsf{all\_tags}:=\emptyset$;
54                                                                            for $l\leftarrow\text{analyzed program points}$:
55                                                                                 for $v\leftarrow\text{struct}\ s\ \text{reachable at}\ l$:
56                                                                                      $N:=\text{possible values of}\ f_{s}\ \text{of}\ v$;
57                                                                                      $\mathsf{all\_tags}:=\mathsf{all\_tags}\cup N$;
58
59
60                                                                                      return $\mathsf{all\_tags}$;
61
Algorithm 2 Identifying tag values associated with fields
The last step of the analysis is to interpret the results of the must-points-to analysis using a heuristic, as demonstrated in Algorithm 2. The entry point is IdentifyTags, which takes three arguments: a candidate struct $s$, a candidate union $u$ in $s$, and a tag-eligible field $f_{s}$ in $s$. Its goal is to determine whether $f_{s}$ serves as a tag field for $u$ and, if it does, to identify the tag values associated with each field of $u$.
As the first step of the heuristic, we invoke CollectFromAccesses (line 1) to identify the associated tag values by examining the value of $f_{s}$ when a union field is accessed after $f_{s}$ has been checked by match/if. This subroutine returns $\mathsf{field\_tags}$, a map from union fields to their tag values, and $\mathsf{all\_tags}$, a set containing all tag values in $\mathsf{field\_tags}$. Initially, both are empty (line 18). We then iterate over every program point in the analyzed functions and check if any field of $u$ is accessed (lines 19–20). If accessed, we determine the possible values of $f_{s}$ from the @$N$ label of the graph computed by the must-points-to analysis (line 21). Here, we treat no-label as the empty set. The possible values should originate from match/if on $f_{s}$, and not from an assignment to $f_{s}$ (line 22). If any of these values are already associated with other fields, we immediately return $\mathsf{None}$, indicating that $f_{s}$ is not a tag field (lines 23–25). Otherwise, we add the values to $\mathsf{field\_tags}$ and $\mathsf{all\_tags}$ (lines 26–27).
If CollectFromAccesses succeeds, IdentifyTags proceeds to the next step by calling CollectFromStructs (line 6). This complements the previous step by discovering tag-field associations that may have been missed due to the analysis’s imprecision. This subroutine considers the field values and the last-written union fields at the end of each basic block. For instance, consider the following code:
(*e).kind = 2; (*e).v.b = ...; return e;
We have the following graph at return:
e @2 .kind.v.b
From this, we can conclude that kind equals 2 when b is the last-written union field, likely associating 2 with b.
We inspect only the states at the end of basic blocks because the states of other program points are prone to provide incorrect information from intermediate states. Consider the following example:
1(*e).kind = 1; (*e).v.e = ...; ...
2(*e).kind = 2;
3(*e).v.b = ...; return e;
Initially, e is used as a negation expression by setting kind to 1, but it becomes an addition expression by setting kind to 2 in the end. If we examine the state after line 2, we will get an incorrect association between 2 and e from the following graph:
e @2 .kind.v.e
To avoid this issue, we examine only the end of basic blocks, where both the tag value and the union field are likely to be correctly set.
However, this approach cannot completely prevent reading intermediate states. For example, a function call can occur between the tag set and the union field set. Therefore, we prioritize the data from CollectFromAccesses, which are less likely to be affected by intermediate states, over CollectFromStructs. If CollectFromAccesses associates a certain tag value with a specific union field, that tag value in CollectFromStructs’ results is ignored.
To implement this approach, CollectFromStructs iterates over the end of every basic block in each analyzed function, identifying every value of type $s$ reachable from local variables at that point (lines 31–33). If the struct’s union has a written field, that field is associated with the values of $f_{s}$ (lines 34–36). Then, IdentifyTags combines results from both subroutines, prioritizing data from CollectFromAccesses (lines 8–13). It removes tags already associated with union fields by CollectFromAccesses from the results of CollectFromStructs (line 9) and returns $\mathsf{None}$ if a tag value remains associated with two fields even after this removal (lines 10–11).
Finally, we search for tag values not associated with any union fields. To achieve this, we call CollectAllTags (line 14), which collects all possible tag values by examining the state at every program point (lines 40–43). By removing the tags associated with union fields, we isolate the tags not associated with union fields (line 15).
To determine the tag field for each union, we run IdentifyTags on all tag-eligible fields. The field for which IdentifyTags returns a value other than $\mathsf{None}$ is identified as the tag field. If multiple tag fields are found, we select the one with the highest number of distinct tag values.
4. Code Transformation
In this section, we present code transformation that replaces unions with tagged unions in C2Rust-generated code, using the results of the static analysis. We demonstrate the transformation of the Expr type defined in Section 2.2 as an example. We assume that the tag field and the tag values associated with each union field are correctly identified by the static analysis.
The type definitions are transformed as follows:
struct Expr { v: C2RustUnnamed }
enum C2RustUnnamed {
Empty0, e1(*mut Expr), b2(BExpr), b3(BExpr) }
struct BExpr { l: *mut Expr, r: *mut Expr }
This result is the same as the hand-written code in Section 2.3, except for the variant names. We generate variant names by concatenating the union field name with the tag value. For tags not associated with any fields, we prepend Empty to the tag value. To improve variant names, one option is to utilize global variables’ names. When using unions with tag fields, C programmers often define an enum, i.e., a group of constant integers with associated names, to use them as tag values. These names typically reflect the programmers’ understanding of the tag values’ meaning, e.g., EXPR_ONE. Since C2Rust translates each enum definition into multiple constant global variable definitions while preserving the names, we can use these variable names instead of the union field names. Although the names generated with this strategy may still be unsatisfactory, developers can easily rename them to more meaningful names using IDEs.
We now discuss the transformation of code using unions. We propose two approaches: naïve transformation, which can be applied to any code but does not adhere to Rust idioms (Section 4.1), and idiomatic transformation, which follows Rust idioms but is applicable only to specific code patterns (Section 4.2). We use both methods within a single codebase, prioritizing idiomatic transformation wherever possible and resorting to naïve transformation when necessary.
4.1. Naïve Transformation
Naïve transformation involves defining helper methods for the transformed structs and unions. These methods are categorized into two groups: reading and writing. We first focus on reading. Below are the read-related methods for Expr and C2RustUnnamed:
1impl Expr {
2 fn kind(self) -> i32 {
3 match self.v {
4 C2RustUnnamed::Empty0 => 0,
5 C2RustUnnamed::e1(_) => 1, ... }}}
6impl C2RustUnnamed {
7 fn get_e(self) -> *mut Expr {
8 if let C2RustUnnamed::e1(v) = self { v }
9 else { panic!() }}}
The kind method (lines 2–5) of Expr replaces the kind field in the original code. This method returns the appropriate tag value by applying pattern matching to the tagged union value. Code that reads a tag field is replaced with a call to the tag-returning method.
The get_e method (lines 7–9) of C2RustUnnamed replaces the union field e. This method returns a value by applying pattern matching to the tagged union value. If the current variant does not contain such a value, the method triggers a panic. This allows the dynamic detection of potential bugs by identifying read from a field other than the last-written one, rather than silently reinterpreting the value. Although not shown in the example, a method get_$f$ is defined similarly for each union field $f$. We replace code reading a union field with a call to the corresponding getter method.
Using these methods, we transform code reading tags and union fields as follows, where the former represents the code before transformation and the latter represents the code after transformation:
match (*e).kind { 1 => eval((*e).v.e), // before
match (*e).kind() { 1 => eval((*e).v.get_e()), // after
Although this transformation preserves the semantics, the resulting code is not idiomatic. It applies pattern matching twice to the tagged union value—once to get the tag value and once to get the union field value—instead of applying pattern matching only once to directly access the associated value of each variant.
We now describe the write-related methods, defined as follows:
1impl Expr {
2 fn set_kind(&mut self, v: i32) {
3 match v {
4 0 => { self.v = C2RustUnnamed::Empty0 }
5 1 => {
6 let v =
7 if let C2RustUnnamed::e1(v) = self.v { v }
8 else { std::ptr::null_mut() };
9 self.v = C2RustUnnamed::e1(v) } ... }}}
10impl C2RustUnnamed {
11 fn deref_e_mut(&mut self) -> *mut *mut Expr {
12 if let C2RustUnnamed::e1(_) = self {}
13 else { *self = Self::e1(std::ptr::null_mut()); }
14 if let C2RustUnnamed::e1(v) = self { v }
15 else { panic!() }}}
The set_kind method (lines 2–9) of Expr replaces assignments to kind. It takes a tag value as an argument and updates the tagged union to the appropriate variant. If the variant has an associated value, we check if this value already exists and reuse it if it does (line 7). If the value does not exist, we create an arbitrary value, which in this case is the null pointer (line 8).
The deref_e_mut method (lines 11–15) of C2RustUnnamed provides a pointer to the inner value, which we use to replace code that mutates e or takes its address. First, we check whether the current variant is appropriate (line 12), and, if not, update the value to the correct variant (line 13). Then, we return a pointer to the value (line 14), while the panic!() on line 15 is never reached. We also define a method deref_$f$_mut for each union field $f$ in a similar manner.
We transform code that updates tags and union fields as follows:
(*e).kind = 1; (*e).v.e = ...; // before
(*e).set_kind(1); *(*e).v.deref_e_mut() = ...; // after
In the transformed code, set_kind changes the variant to e1. Then, deref_e_mut returns a pointer to the arbitrary value set by set_kind, and the indirect assignment to the pointer updates this value.
Note that deref_e_mut does not trigger a panic even when the current variant differs from the expected one, unlike get_e. This behavior ensures the correct transformation of code that first writes to a union field and then sets the tag. For example, the following transformation preserves the semantics:
(*e).v.e = ...; (*e).kind = 1; // before
*(*e).v.deref_e_mut() = ...; (*e).set_kind(1); // after
After the transformation, deref_e_mut changes the variant to e1, and the indirect assignment to the pointer sets the associated value. Then, set_kind retains both the variant and associated value. Although this approach preserves the semantics, it is not idiomatic in Rust, as we can create a tagged union value within a single expression instead of setting the tag and the union field separately.
4.2. Idiomatic Transformation
Idiomatic transformation uses pattern matching on tagged union values and constructs a tagged union value with a single expression, avoiding helper methods. Below is the transformation of match:
match (*e).kind { 1 => eval((*e).v.e), // before
// after
match (*e).v { C2RustUnnamed::e1(ref x) => eval(*x),
The match condition is the tagged union value itself, rather than the tag value obtained by get_kind. In addition, each branch directly matches the variant, instead of comparing the tag value to an integer. The ref keyword before the identifier x binds a pointer to the associated value, not the value itself, to x, allowing the branch to read and modify the value. This eliminates the need to call methods such as get_e and deref_e_mut, as x directly accesses the value.
We also transform if that checks tag values to use pattern matching with if-let, as illustrated in the following example:
if (*e).kind == 1 { eval((*e).v.e) } // before
// after
if let C2RustUnnamed::e1(ref x) = (*e).v { eval(*x) }
We handle the disjunction of multiple equality comparisons using |, the or operator in pattern matching. An example is shown below:
if (*e).kind == 2 || (*e).kind == 3 { eval((*e).v.b.l) }
// before
if let C2RustUnnamed::b2(ref x) | // after
C2RustUnnamed::b3(ref x) = (*e).v { eval((*x).l) }
However, the idiomatic approach is not applicable to other kinds of conditions, such as conjunctions and inequalities, because Rust’s patterns currently cannot represent these conditions. A conjunction specifies that the tag is a particular integer and also that some boolean formula is satisfied. When using match for pattern matching, Rust offers match guards (Rust, 2024f) to express such logic. However, when using if-let, this logic is supported through let chains (Farrokhzad, 2018), which is currently an unstable feature and requires a special flag to enable. For this reason, we chose not to transform conjunctions into pattern matching in this work, leaving it for future work once this feature is stabilized. On the other hand, an inequality specifies that the tag is not a particular integer. However, the purpose of pattern matching is to check whether a value conforms to a certain pattern, not that it does not, and thus it cannot replace inequalities. Thus, the following code is transformed using the naïve approach:
if (*e).kind == 1 && ... { eval((*e).v.e) }
if (*e).kind != 1 { ... } else { eval((*e).v.e) }
The idiomatic transformation is also not applicable when union fields are accessed without using match or if to check the tag field. C programmers sometimes assume that a tag field has a specific value at a certain point based on their understanding of the program’s behavior and directly access the union field without checking the tag. Since tag check is not performed, we cannot transform such code into pattern matching and must resort to the naïve transformation.
The idiomatic transformation consolidates multiple assignment expressions within a single code block into a single assignment expression that constructs a tagged union value if the assignments set the tag and union fields. Below illustrates this transformation:
{ ... (*e).kind = 1; (*e).v.e = ...; ... } // before
{ ... (*e).v = C2RustUnnamed::e1(...); ... } // after
When multiple assignments are distributed across different code blocks, they are individually transformed using the naïve approach.
5. Evaluation
In this section, we evaluate our approach with 36 real-world C programs. We first describe our implementation of Urcrat, which realizes the proposed approach (Section 5.1), and the process of collecting the benchmark programs (Section 5.2). The implementation and benchmark programs are publicly available (Hong and Ryu, 2024). We then assess our approach by addressing the following research questions:
  • RQ1. Precision and recall: Does it identify tag fields without false positives or false negatives? (Section 5.3)
  • RQ2. Correctness: Does it transform code while preserving its semantics? (Section 5.4)
  • RQ3. Efficiency: Does it efficiently analyze and transform programs? (Section 5.5)
  • RQ4. Code characteristics: How much does the code change due to the transformation, and how frequently are the helper methods called? (Section 5.6)
  • RQ5. Impact on performance: What is the effect of replacing unions with tagged unions on program performance? (Section 5.7)
Our experiments were conducted on an Ubuntu machine with Intel Core i7-6700K (4 cores, 8 threads, 4GHz) and 32GB DRAM. Finally, we discuss potential threats to validity (Section 5.8).
5.1. Implementation
We built Urcrat on top of the Rust compiler (Rust, 2024a). Urcrat analyzes Rust code after it has been lowered to Rust’s mid-level intermediate representation (MIR) (Rust, 2024c), which expresses functions as control flow graphs with basic blocks. For code transformation, Urcrat utilizes Rust’s high-level intermediate representation (HIR) (Rust, 2024b), akin to abstract syntax trees but with syntactic sugar removed and symbols resolved. We employed C2Rust v0.18.0 with minor modifications.
5.2. Benchmark Program Collection
Table 1. Benchmark programs
Name C LOC Rust LOC #Unions #Candidates #Identified
bc-1.07.1 10810 16982 4 1 1
binn-3.0∗∗ 5686 4298 1 1 0
brotli-1.0.9∗∗ 13173 127691 6 4 0
cflow-1.7 20601 26375 5 4 3
compton 8748 14084 2 2 2
cpio-2.14 35934 80929 10 4 3
diffutils-3.10 59377 95835 7 5 4
enscript-1.6.6 34868 78749 9 5 3
findutils-4.9.0 80015 139858 13 6 3
gawk-5.2.2 58111 140566 17 10 3
glpk-5.0 71805 145738 18 14 3
gprolog-1.5.0 52193 74381 5 2 0
grep-3.11 64084 84902 11 9 6
gsl-2.7.1 227199 422854 14 14 0
gzip-1.12 20875 21605 4 2 1
hiredis 7305 14042 1 1 1
make-4.4.1 28911 36336 1 1 1
minilisp 722 2149 1 1 1
mtools-4.0.43 18266 37021 2 1 0
nano-7.2 42999 74994 6 4 3
nettle-3.9 61835 82742 5 2 1
patch-2.7.6 28215 103839 3 1 1
php-rdkafka 3771 28864 1 1 1
pocketlang 14267 41439 4 3 3
pth-2.0.7 7590 12950 1 1 1
raygui 1588 17218 1 1 1
rcs-5.10.1 28286 36267 1 1 1
screen-4.9.0 39335 72199 1 1 0
sed-4.9 48190 68465 8 7 4
shairport 4995 10118 2 1 1
tar-1.34 66172 134972 16 12 9
tinyproxy 5667 12825 5 2 2
twemproxy 22738 74593 8 7 5
uucp-1.07 51123 77872 3 3 0
webdis 14369 29474 2 2 2
wget-1.21.4 81188 192742 6 5 4
Total 204 141 74
: from GitHub, ∗∗: from Crown
We collected benchmark programs from three sources: (1) Crown (Zhang et al., 2023), (2) GNU packages (GNU, 2024), and (3) GitHub. Only 2 out of 20 programs used by Crown have candidate unions, so we expanded the benchmarks with additional sources. We chose GNU packages for their representative C projects and GitHub for its diverse code patterns. We avoided large codebases because C2Rust often produces Rust code with type errors, primarily due to missing type casts, which require significant manual effort to correct. From GNU, we gathered C packages with less than 250k LOC and individual Wikipedia entries, indicating they are well-known. From GitHub, we gathered C projects with less than 1 MB of code and over 1,000 stars. In both collections, we retained programs (1) compiled successfully on Ubuntu, (2) transpiled successfully by C2Rust, and (3) containing candidate unions. This resulted in 24 programs from GNU and 10 from GitHub, giving us a total of 36 benchmark programs. Of these, 20 required manual fixes after C2Rust’s translation, with an average of 26.6 lines modified. Columns 2–5 of Table 1 present the C LOC, Rust LOC, number of unions, and number of candidate unions in each program, respectively. In the benchmarks, 21% of unsafe functions involve unions, and 11% involve tagged unions.
5.3. RQ1: Precision and recall
We evaluate the precision and recall of the proposed approach. We first identified the tag field of each candidate union through static analysis. Column 6 of Table 1 shows the number of candidate unions for which a tag field is identified in each benchmark program. Out of 141 candidates across 36 programs, Urcrat identifies tag fields for 74 candidates in 29 programs. We then manually inspected each candidate union to determine the presence of a tag field, checking for false positives (a tag field identified for a union that does not have one) and false negatives (no tag field identified for a union that has one). For practical use of our tool, false positives are problematic because they change the program’s semantics. Conversely, false negatives are less problematic as they only prevent the replacement of unions with tagged unions, and the semantics remains correct.
Our manual inspection shows that the static analysis is precise, revealing no false positives and only five false negatives. This results in a precision of 100% and a recall of 93.7% ($=74/79$). Specifically, two false negatives occur in glpk-5.0, while the others are in gawk-5.2.2, screen-4.9.0, and uucp-1.07, respectively. Three of these (from glpk-5.0, gawk-5.2.2, and screen-4.9.0) are due to intermediate states involving tag values that are not associated with any union fields by CollectFromAccesses. We now discuss the reasons for the remaining two false negatives:
glpk-5.0
The false negative arises from goto in the C code:
switch (tab->type) { case 112: goto input_table;
case 119: goto output_table; default: abort(); }
input_table: ... return; output_table: ...
In this code, type is the tag field. Since Rust does not support goto, C2Rust translates the code as follows:
match (*tab).type {
112 => current_block = 1, 119 => { ... } _ => abort(),}
if current_block == 1 { ... }
The resulting code has a variable, current_block, which mimics goto’s effect. As the analysis is path-insensitive, it concludes that both 112 and 119 are possible values in the true branch of if, hindering the identification of the tag field. To address this issue, we need to either analyze the original C code or modify the C code to avoid goto. We chose the latter and revised the code as follows:
switch (tab->type) { case 112: ... break;
case 119: ... break; default: abort(); }
Urcrat can identify the tag field from the modified version.
uucp-1.07
This false negative is due to a bug in the C code:
if (qport->uuconf_ttype == 5) {
if (qport->uuconf_u.uuconf_stli.zdevice != NULL)
fprintf(e,"%s",qport->uuconf_u.uuconf_smodem.zdevice);
Here, uuconf_ttype is the tag field, and the union field associated with 5 is uuconf_stli. However, the code erroneously accesses uuconf_smodem in the fprintf statement. The function contains multiple fprintf statements, suggesting that this bug was likely introduced through copy-pasting. After correcting the code to access uuconf_stli, Urcrat successfully identifies the tag field.
5.4. RQ2: Correctness
We evaluate the correctness of our approach by checking whether the transformed program is compilable and exhibits the same behavior as the original. For the experiments, we used the fixed code for glpk-5.0 and uucp-1.07, allowing the identification of additional tag fields. Consequently, we examined 30 programs: 29 identified without code fixes and 1, which is uucp-1.07, identified only after the code fix. All 30 programs are compilable after transformation. Among 23 programs with test suites, 17 pass the tests post-transformation. Those failed are gawk-5.2.2, grep-3.11, make-4.4.1, minilisp, twemproxy, and wget-1.21.4.
We manually investigated the reasons for the failures and found that only gawk-5.2.2 and twemproxy’s failures result from imprecise identification of tag values in the static analysis. The other failures are due to two specific C code patterns requiring minor manual code fixes after transformation: (1) reading a union field other than the last-written one, and (2) code relying on memory layout. We now discuss the failure reasons for each program.
gawk-5.2.2
The failure arises from tag values not being identified due to intraprocedural analysis.
INSTRUCTION *bcalloc(int op, ...) {
INSTRUCTION *cp = ...; cp->opcode = op; ... return cp;}
In this code, opcode is the tag field, assigned the argument value. The analysis cannot identify values at the call-site of bcalloc as possible opcode values. This causes the transformed program to panic when an unknown tag value is passed to set_opcode.
twemproxy
The failure is due to the lack of C library modeling, preventing identification of tag values.
getaddrinfo(n, s, &h, &ai); si->family = ai->ai_family;
Here, family is the tag field, and ai_family determines its value, set by the libc function getaddrinfo. While 2 and 10 are possible values, the analysis fails to recognize them as tag values. This issue can be resolved by incorporating library modeling into the analysis.
grep-3.11
It fails because of reading a field not lastly written.
fetch_token(token, input, syntax);
c = token->opr.c; if (token->type == 2) return;
In this code, type is the tag field, and opr is the union value. The code first reads c and then checks type, causing a panic when get_c is called in the transformed code. Swapping the order of these statements allows the transformed program to pass the tests.
make-4.4.1
This is also due to reading a field not lastly written.
struct function_table_entry { ..., int alloc_fn;
union {fptr1 func_ptr; fptr2 alloc_func_ptr;} fptr;};
if (!entry_p->fptr.func_ptr) abort();
The union has two fields, func_ptr and alloc_func_ptr, with alloc_fn as the tag field. The code checks if func_ptr is null regardless of alloc_fn’s value, exploiting that different function pointer types have the same size and null representation. After transformation, this causes a panic if alloc_func_ptr is the last-written field. The code can be fixed as follows to pass the tests post-transformation:
if (entry_p->alloc_fn == 0 && !entry_p->fptr.func_ptr ||
entry_p->alloc_fn==1 && !entry_p->fptr.alloc_func_ptr)
minilisp
This failure is caused by different memory layouts of unions and tagged unions.
struct Obj { int type; int size; union { ... }; };
Obj *alloc(size_t size, ...) { size += 8; ... }
The alloc function allocates an Obj in a global byte array. It determines the memory size by taking the size of a union field and adding 8 to account for the offset of the union within Obj. After transformation, the code becomes as follows:
struct Obj {size: i32, c2rust_unnamed: C2RustUnnamed_1}
Now, each variant value is at offset 16 due to alignment requirements, resulting in alloc allocating insufficient memory. We corrected the code to allocate larger memory, enabling the transformed program to pass the tests.
wget-1.21.4
It also stems from different memory layouts.
addr->family = 2; memcpy(&addr->data, tmp, 4);
Here, family is the tag field, and 2 is associated with the union field d4 of addr->data. The C program writes to addr->data because addr->data and addr->data.d4 denote the same address. However, after transformation, memcpy overwrites the tag at offset 0 of addr->data. We could pass the tests by fixing the code as follows, facilitating deref_d4_mut to be called in the transformed code:
addr->family = 2; memcpy(&addr->data.d4, tmp, 4);
Currently, for end users of the tool, detecting and fixing incorrect translations is challenging. To identify incorrect translations, they must manually inspect the code or run tests. If test suites do not exist, creating new test cases can be costly. Additionally, test cases may not always reveal incorrect translations. Fixing incorrect code is even more challenging, as users need to identify the root cause. The difficulty of this process depends on the root cause. When tag values are not correctly identified or fields not lastly written are read, investigating the cause is relatively straightforward. The test fails due to a panic, allowing users to check which tag value or field read triggered the panic. However, issues related to different memory layouts are more difficult to diagnose since they do not trigger panics. In such cases, users must rely on their debugging skills. We believe it would be beneficial to identify common incorrect translation patterns and design analyses to detect these patterns, thereby providing users with warnings. We leave this as future work.
5.5. RQ3: Efficiency
Figure 2. Urcrat execution time
We evaluate the efficiency of the proposed approach by measuring the execution time of Urcrat, which includes both analysis and transformation, for each program. Fig. 2 presents the execution time relative to the Rust LOC, with the y-axis displayed in a log scale due to the wide range of execution times. Urcrat efficiently handles most programs, with 31 programs taking less than a minute. The longest execution time is 4,910 seconds for gawk-5.2.2. Execution time shows only a weak correlation with code size, as other factors, primarily the complexity of pointer graphs and the number of analyzed functions, also significantly influence execution time.
We also investigate the effectiveness of our selective function analysis in terms of efficiency. For each program, the transformation consumes less than 1% of the total time, as it involves only tree-walking, while the analysis occupies the remaining time. This highlights the importance of reducing analysis time. We calculated the percentage of analyzed functions relative to the total functions in each program, finding a geometric mean of 4.80%. In 28 programs, the analyzed functions constitute less than 10% of the total functions. The highest percentage is 58.23% in the case of minilisp. These results indicate that our approach is effective for efficiency.
5.6. RQ4: Code Characteristics
To understand the characteristics of the transformed code, we first evaluate the extent of code changes introduced. We measured the changes in the 30 programs where tag fields are identified. On average, the transformation added 861.8 lines per program due to helper method definitions. Excluding these, an average of 252.4 lines were inserted and 301.8 lines were deleted per program. These results demonstrate the practical utility of our approach, as manually implementing such significant code modifications would be both time-consuming and error-prone.
We also evaluate the applicability of the idiomatic transformation, which avoids inserting method calls. With only the naïve transformation, each program has 194.6 method calls on average. In contrast, applying the idiomatic transformation alongside the naïve transformation reduces this to 124.8 method calls per program, achieving a 36% decrease. Specifically, 38.3 calls are removed by replacing match on tag values with match on tagged unions, 15.8 calls by replacing if with if-let, and 15.6 calls by consolidating separate assignments into tagged union construction. The results indicate that the idiomatic transformation improves code quality.
5.7. RQ5: Impact on Performance
We evaluate the impact of replacing unions with tagged unions on the performance of the translated programs. We compare the performance of each Rust program before and after the transformation, measured in terms of the execution time of the test suite. To ensure reliable results, we excluded test suites with execution times shorter than 0.1 seconds, resulting in 20 programs, and computed the average execution time from fifty runs for each program.
The experimental results show that the performance overhead is negligible. On average, the transformed programs were only 0.01% slower than the original ones. Among the 20 programs, 5 ran slightly slower after transformation, while the others were faster. We also performed one-sided t-test with the null hypothesis that the program after transformation is slower than the original program by at least 1%. With a significance level of 0.05, we could reject the null hypothesis for 9 out of 20 programs, implying that the transformation is unlikely to incur observable performance overhead for them. For the remaining 11 programs, we could not reject the null hypothesis, but this does not necessarily mean they have an overhead greater than 1%. Repeating the experiments may provide stronger statistical significance, rejecting the null hypothesis.
5.8. Threats to Validity
Threats to external validity primarily stem from the choice of benchmarks. GNU packages often share common code patterns, which may introduce bias. Although we included projects from GitHub to enhance diversity, this may not fully represent the entire C ecosystem. Notably, large codebases might exhibit different characteristics. Further experiments with a broader range of C programs would provide greater confidence in the generalizability of our approach.
The type errors produced by C2Rust also pose a threat to external validity. Before using our tool, all type errors in C2Rust-generated code must be resolved. Larger programs tend to exhibit more type errors after C2Rust’s translation, requiring significant effort to correct. This could discourage users from adopting both C2Rust and Urcrat, potentially impeding the widespread adoption of our approach. To mitigate this issue, we can improve C2Rust or develop techniques for fixing type errors in C2Rust-generated code.
Threats to construct validity arise from using test suites for correctness and performance evaluation. Passing all tests does not guarantee correctness. However, tests are the most widely used method for practical semantics validation and successfully discovered incorrect behavior in some programs during our evaluation. For performance assessment, the used test suites may be inadequate as they were not designed for performance measurement.
6. Related Work
Improving C2Rust-Generated Code
Several studies have proposed techniques to improve C2Rust’s translation by replacing unsafe features with safe counterparts in Rust. However, none of these studies address the translation of unions, which is the focus of this work. Laertes (Emre et al., 2021, 2023) and Crown (Zhang et al., 2023) aim to replace raw pointers with references. Laertes uses compiler feedback to determine which pointers to replace and the appropriate lifetimes for the references. In contrast, Crown employs ownership analysis to facilitate the replacement of a larger number of pointers compared to Laertes. Concrat (Hong and Ryu, 2023) targets the replacement of some external function calls with equivalent functions from the Rust standard library. It employs dataflow analysis to identify the use of locks, allowing for the substitution of the C lock API with the Rust lock API.
C-to-Rust Translation with Language Models
Recent studies have developed language models capable of translating code (Roziere et al., 2020; Lachaux et al., 2024; Rozière et al., 2022; Szafraniec et al., 2023; Liu et al., 2023; Wang et al., 2021; Chen et al., 2021; Feng et al., 2020; Guo et al., 2021; Touvron et al., 2023) or proposed techniques to utilize pre-trained language models for code translation (Yang et al., 2024; Pan et al., 2024). Most focus on translating between languages like C++, Java, and Python, rather than C to Rust. A notable exception is Rozière et al. (2022), which translates C++ to Rust, but their evaluation uses only small competitive programming solutions that do not involve unions, with only 21% translated correctly. Similarly, Pan et al. (2024) use GPT-4 for C-to-Rust translation, applying it to CodeNet (Puri et al., 2021), which consists of programming problem solutions, achieving 61% correctness. These results indicate that language models frequently generate code that is uncompilable or semantically incorrect, while our approach consistently produces compilable and mostly correct code.
Must-Points-To Analysis
Researchers have studied must-points-to analysis, but their techniques differ from ours in purpose, method, and target language. Our analysis aims to precisely compute struct field values. In contrast, most studies use must-points-to analysis to enhance the precision of other analyses. Altucher and Landi (Altucher and Landi, 1995) use must-points-to relations to compute def-use relations. Ma et al. (Ma et al., 2008) and Fink et al. (Fink et al., 2008) improve may-points-to relations using must-points-to relations, enabling more precise detection of null pointer dereferences and Java typestate checking, respectively. Nikolić and Spoto (Nikolić and Spoto, 2012) use must-points-to analysis to improve other analyses, such as nullness and termination analyses. While our approach uses may-points-to analysis as a pre-analysis, some techniques compute may- and must-points-to relations simultaneously. Emami et al. (Emami et al., 1994) conduct interprocedural may- and must-points-to analysis for compiler optimizations and parallelizations. Sagiv et al. (Sagiv et al., 1999) propose a parametric framework that generates a family of shape analyses based on three-valued logic, expressing must-, must-not-, and may-points-to relations. Unlike most studies, including ours, which focus on imperative languages, Jagannathan et al. (Jagannathan et al., 1998) propose must-points-to analysis for functional languages, with results used for optimizations such as closure conversion. Techniques applicable to general must-points-to analysis have also been explored. Balatsouras et al. (Balatsouras et al., 2017) present a declarative model in Datalog that expresses a wide range of must-points-to analyses. Kastrinis et al. (Kastrinis et al., 2018) propose an efficient data structure for pointer graphs.
7. Conclusion
In this work, we address the challenge of transforming unions with tags into tagged unions in C2Rust-generated code. This necessitates identifying tag fields for unions and the tag values associated with union fields. To achieve this, we propose a static analysis method that includes must-points-to analysis for computing struct field values and a heuristic for interpreting this information. In addition, we present a code transformation technique that generates idiomatic code for specific patterns while ensuring semantics preservation for the remaining code. Our evaluation shows that the proposed approach is precise, mostly correct, and scalable.
Acknowledgements.
This research was supported by the National Research Foundation of Korea (NRF) (2022R1A2C200366011 and 2021R1A5A1021944), the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (2022-0-00460, 2023-2020-0-01819, and 2024-00337703), and Samsung Electronics Co., Ltd (G01210570).